# Convex Geometry of Building Sets

### v3

@@ Abstract.
Building sets were introduced in the study of wonderful compactifications of hyperplane arrangement complements and were later generalized to finite meet-semilattices. Convex geometries, the duals of antimatroids, offer a robust combinatorial abstraction of convexity. Supersolvable convex geometries and antimatroids appear in the study of poset closure operators, Coxeter groups, and matroid activities. We prove that the building sets on a finite meet-semilattice form a supersolvable convex geometry. As an application, we demonstrate that building sets and nested set complexes respect certain restrictions of finite meet-semilattices unifying and extending results of several authors.
@@ 1. Introduction
De Concini and Procesi’s celebrated wonderful compactifications of a hyperplane arrangement complement proceed by blow ups of certain flats of the arrangement [14]. There are choices for which flats are blown up, and each set of choices is encoded by a building set. Feichtner and Kozlov [20] extended the notion of a building set from the lattice of flats of a hyperplane arrangement to arbitrary finite meet-semilattices [^1] Henceforth, all posets in this paper will be assumed to be finite., and observed that this generalization admits connections to other topics such as stellar subdivisions of fans.
@@
The Bergman fan of a matroid was introduced by Ardila-Klivans [5] as a matroidal generalization of the tropicalization of a linear space. Feichtner-Sturmfels [21] showed that the nested set complex associated to a building set for the lattice of flats gives a unimodular triangulation of the Bergman fan. In seminal work of Adiprasito-Huh-Katz [1], the Chow ring of a matroid, introduced earlier by Feichtner-Yuzvinsky [22], was utilized for settling the Heron-Rota-Welsh conjecture on the log-concavity of the coefficients of the characteristic polynomial of a matroid. Although Feichtner-Yuzvinsky’s construction is defined for arbitrary building sets, Adiprasito-Huh-Katz found that the maximum building set was sufficient for their goals. [^2] While the Bergman fans of [1] are not Bergman fans induced by building sets on the lattice of flats, we note a connection in footnote 3.17. Much of the subsequent work by researchers studying Chow rings of matroids has been focused on the maximum building set with a few important notable exceptions: the conormal fan of a matroid [^3] The biflats, viewed as a subset of $\mathcal{L}(M)^{op}\times\mathcal{L}(M^{\perp})$, are closed under taking joins and thus satisfy condition (2) of Proposition 2.11. While the biflats do not satisfy condition (1) of Proposition 2.11, the image of the biflats in $\Sigma_{M}\times\Sigma_{M^{\perp}}$ contains the image of the irreducibles. As the biflats are closed under taking joins, the would-be nested set complex associated to the biflats is the order complex, of which the conormal fan is a distinguished codimension-2 subcomplex. This observation will be elaborated upon in forthcoming work of the first author and Castillo. [3, 4, 30], the augmented Bergman fan of a matroid [10, 11, 18, 27, 29], and the Bergman fan of a polymatroid [13, 19, 31]. [^4] It follows from each of the works [2, 3, 31] that the Kähler package holds for the Chow ring of a matroid with respect to an arbitrary building set.
@@
Perhaps surprisingly, there has been no work exploring the systematic construction of building sets for the lattice of flats of a matroid – this apparent gap in the literature was the genesis of this paper. Although our primary motivation is the study of matroids, we have found that our results hold for arbitrary meet-semilattices. In their original article, de Concini-Procesi provided three different equivalent descriptions of a building set [14]. Their most elegant characterization was popularized by Postnikov [32] and Feichtner-Sturfmels [21] for the Boolean lattice, but has largely been overlooked by researchers studying matroids. We begin by extending this characterization to a general meet-semilattice $\mathcal{L}$, e.g. a geometric lattice. We apply this result to demonstrate that the building sets for $\mathcal{L}$ form an intersection lattice. It follows that given an arbitrary subset $X\subseteq\mathcal{L}$, there exists a smallest building set $B$ which contains $X$, thus the building sets for $\mathcal{L}$ determine a closure operator; this generalizes a result of Feichtner-Sturmfels [21] from the Boolean lattice to general meet-semilattices.
@@
Convex geometries are a special class of closure systems introduced independently by Edelman and Jamison [16, 17, 25] providing a combinatorial abstraction of convex hulls in finite point configurations and upper ideals in posets. A fundamental result of Björner [8] states that convex geometries are dual to antimatroids, which are an important class of greedoids. Supersolvable antimatroids and convex geometries were introduced implicitly in the work of Stanley [34], and were developed in earnest by Armstrong [6]. They arise naturally in the study of poset closure operators [24], Coxeter groups [6], and matroid activities [23]. We prove that the building sets on a meet-semilattice form a supersolvable convex geometry.
@@
We demonstrate the utility of our results by proving that the building sets and nested set complexes behave well with respect to restrictions of meet-semilattices. This result unifies the work of Mantovani-Padrol-Pilaud on oriented matroids [28] with Bergman fan constructions arising in the study of Hodge theory for matroids due to Braden-Huh-Matherne-Proudfoot-Wang [11], Crowley-Huh-Larson-Simpson-Wang [13], and Eur-Larson[19].
@@ 2. Semilattices and Building Sets
In this section we review the theory of posets, semilattices, and building sets. We prove Proposition 2.11 which extends a characterization of building sets due to de Concini and Procesi from intersection lattices of subspace arrangements to general meet-semilattices. The section concludes with Proposition 2.12 which states that the collection of building sets on a meet-semilattice ordered by inclusion forms an intersection lattice generalizing an observation of Feictner-Sturmfels from the Boolean lattice to general meet-semilattices.
@@
We write $x<y$ if $x\leq y$ and $x\neq y$, and we define $x>y$ similarly. We say that $y$ covers $x$ if $x<y$ and there is no element $z\in\mathcal{P}$ such that $x<y<z$. We define the subset $\mathcal{P}_{\leq x}$ to be $\{z\in\mathcal{P}:z\leq x\}$. For a subset $B$ of $\mathcal{P}$ with the induced order, and $x\in\mathcal{P}$, we define $B_{\leq x}=\{z\in B:z\leq x\}$, and $\text{max}\,B_{\leq x}$ as the set of maximal elements in $B_{\leq x}$ with respect to $\leq$. An interval $[x,y]\subseteq\mathcal{P}$ is the set of elements $z\in\mathcal{P}$ such that $x\leq z\leq y$. A linear extension $<_{L}$ of a partially ordered set $(\mathcal{P},\leq)$ is a total order of $\mathcal{P}$ such that if $x,y\in\mathcal{P}$ with $x<y$, then $x<_{L}y$. Let $(\mathcal{P}_{1},\leq_{1})$ and $(P_{2},\leq_{2})$ be posets. A function $f:\mathcal{P}_{1}\rightarrow\mathcal{P}_{2}$ is an order embedding if $x\leq_{1}y$ precisely when $f(x)\leq_{2}f(y)$. If $f$ is the identify map, we may write $\mathcal{P}_{1}\hookrightarrow\mathcal{P}_{2}$. An order embedding is an isomorphism if it is bijective. Let $\{(\mathcal{P}_{k},\leq_{k}):1\leq k\leq n\}$ be a collection of posets. The product $\prod_{k=1}^{n}\mathcal{P}_{k}$ is naturally equipped with a partial order $\leq$ where for $x,y\in\prod_{k=1}^{n}\mathcal{P}_{k}$ we have $x\leq y$ if and only if $x_{k}\leq_{k}y_{k}$ for all $k$ with $1\leq k\leq n$. If a poset has a minimum element, we may denote this element by $\hat{0}$, and if a poset has a maximum element, we may denote this element by $\hat{1}$. An element $a\in\mathcal{P}$ is an atom if it covers $\hat{0}$. A poset is irreducible if it is not a direct product of two other posets, each consisting of at least two elements. Let $I(\mathcal{P})=\{x\in\mathcal{P}\mid[\hat{0},x]\leavevmode\nobreak\ \text{is irreducible}\}$ be the set of irreducibles of $\mathcal{P}$. The set $\text{max}\,I(\mathcal{P})_{\leq x}$ is the set of elementary divisors of $x$.
@@
[20, Proposition 2.1] Let $(\mathcal{P},\leq)$ be a partially ordered set with a minimum element $\hat{0}$ and a maximum element $x$. Let $\text{max}\,I(\mathcal{P})=\{y_{1},...,y_{l}\}$, then
@@
If $x$ is irreducible, then the statement is trivial. Otherwise $x$ is reducible and, by induction, there exists some factorization $[\hat{0},x]=\mathcal{P}\cong\prod_{i=1}^{k}\mathcal{P}_{i}$ with $k\geq 2$, such that $\mathcal{P}_{i}$ is irreducible and $|\mathcal{P}_{i}|\geq 2$ for each $i$. We claim that, after some potential reordering, $\mathcal{P}_{i}\cong[\hat{0},y_{i}]$ and $k=l$. First we observe that each $\mathcal{P}_{i}$ must have a minimum element, which we denote $\hat{0}$, and a maximum element $z_{i}$, otherwise $\mathcal{P}$ would have more than one minimal or maximal element. Let $\mathcal{P}_{i}=[\hat{0},z_{i}]$ and let $\phi:\prod_{i=1}^{k}\mathcal{P}_{i}\rightarrow\mathcal{P}$ be an explicit isomorphism. For any fixed $y_{j}$, there exists $w_{i}\in\mathcal{P}_{i}$ for $1\leq i\leq k$ such that $\phi(\prod_{i=1}^{k}w_{i})=y_{j}$. It follows from the definition of a poset isomorphism that $[\hat{0},y_{j}]\cong\prod_{i=1}^{k}[\hat{0},w_{i}].$ Therefore, there is a unique index $i$ for which $w_{i}\neq\hat{0}$. By assumption, each $z_{i}$ is irreducible. Letting $\phi(z_{i})$ denote $\phi(\hat{0},\ldots,z_{i},\ldots,\hat{0})$, we have $y_{j}\leq\phi(z_{i})$. By the definition of $y_{j}$, it must be that $y_{j}=\phi(z_{i})$. Thus the map $\phi$ induces a canonical bijection between $\{z_{1},\ldots,z_{k}\}$ and $\{y_{1},\ldots,y_{l}\}$, and the statement of the claim follows. ∎
@@
Let $\mathcal{P}$ be a poset and $x,y\in\mathcal{P}$. The greatest lower bound of elements $x$ and $y$, if it exists, is the meet of $x$ and $y$, denoted $x\wedge y$. The least upper bound of $x$ and $y$, if it exists, is the join of $x$ and $y$, denoted $x\vee y$. A poset $\mathcal{L}$ is called a meet-semilattice if the meet of any two of its elements exists in $\mathcal{L}$. We define a join-semilattice similarly. A poset $\mathcal{L}$ is called a lattice if it is a meet-semilattice and a join-semilattice. For avoiding confusion we may denote $\wedge$ as $\wedge^{\mathcal{L}}$, and $\vee$ as $\vee^{\mathcal{L}}$. If $\mathcal{L}$ is a meet-semilattice, we define +:=\$\mathcal{L}^{+}:=\mathcal{L}\backslash\{\hat{0}\}$. Let $\mathcal{L}$ and $\mathcal{K}$ be meet-semilattices. A meet-semilattice order embedding $f:\mathcal{L}\rightarrow\mathcal{K}$ is an order embedding of $\mathcal{L}$ and $\mathcal{K}$ as posets such that for every $x,y\in\mathcal{L}$, we have $x\wedge^{\mathcal{L}}y=f(x)\wedge^{\mathcal{K}}f(y)$.
@@
Thus if $\mathcal{L}$ is a meet-semilattice, or a join-semilattice, and $x,y\in\mathcal{L}$ with $x\leq y$, the interval $[x,y]\subseteq\mathcal{L}$ is itself a lattice. Let $\mathcal{E}$ be a set and $\mathbb{S}$ a lattice of subsets of $\mathcal{E}$ with meet given by intersection, then we refer to $\mathbb{S}$ as an intersection lattice. [^5] We warn the reader that this terminology, although natural, could be a source of confusion as the lattice of flats of a hyperplane arrangement form an intersection lattice, but this is the reverse of the intersection order of flats in an arrangement. In de Concini-Procesi [14], this confusion is avoided by working in the dual vector space.
@@
Let $\mathcal{L},\mathcal{K}$ be meet-semilattices, $f:\mathcal{L}\rightarrow\mathcal{K}$ a meet-semilattice order embedding, and $x_{1},\ldots,x_{k}\in\mathcal{L}$ such that
$\bigvee_{1\leq i\leq k}^{\mathcal{L}}x_{i}\in\mathcal{L}\,,\,\,\,\,\text{then}\,\,\,\,\bigvee_{1\leq i\leq k}^{\mathcal{K}}x_{i}\leq\bigvee_{1\leq i\leq k}^{\mathcal{L}}x_{i}\,.$
@@
Because $f$ is an order embedding, we can identify $\wedge^{\mathcal{K}}=\wedge^{\mathcal{L}}$ and denote this meet by $\wedge$. Let $\mathcal{S}=\{y\in\mathcal{K}:\exists\,i\in[k],\,x_{i}\leq y\}$ and $\mathcal{T}=\{y\in\mathcal{L}:\exists\,i\in[k],\,x_{i}\leq y\}$, then $\mathcal{S}\supseteq\mathcal{T}$. Hence
@@
[14, Section 2.3][20, Definition 2.2] Let $\mathcal{L}$ be a meet-semilattice, and $B\subseteq\mathcal{L}^{+}$. The set $B$ is a building set of $\mathcal{L}$ if for any $x\in\mathcal{L}^{+}$ and $\text{max}\,B_{\leq x}=\{x_{1},...,x_{k}\}$ there is an isomorphism of posets
@@
with $\phi_{x}(y_{1},...,y_{k})=y_{1}\vee\cdots\vee y_{k}$ for $y_{j}\in[\hat{0},x_{j}]$ with $1\leq j\leq k.$ We call $\text{max}\,B_{\leq x}$ the set of factors of $x$ in $B$.
@@
Let $\mathcal{L}$ be a meet-semilattice and $B\in\mathbb{B}(\mathcal{L})$. A subset $N\subseteq B$ is a nested set if for every $x_{1},...,x_{t}\in N$ with $t\geq 2$, which are incomparable elements, $\𝐵\bigvee_{i=1}^{t}x_{i}\in\mathcal{L}\backslash B$. The nested set complex $\mathcal{N}(B)$ is the abstract simplicial complex consisting of all nested sets of $B$. For avoiding confusion, we may denote $\mathcal{N}(B)$ as $\mathcal{N}_{\mathcal{L}}(B)$.
@@
[14, Section 2.4][20, Proposition 2.8] Let $\mathcal{L}$ be a meet-semilattice, $B\in\mathbb{B}(\mathcal{L})$, and $\mathcal{N}(B)$ the associated nested set complex. Let $N\in\mathcal{N}(B)$ with $x_{1}\ldots,x_{k}\in N$ incomparable and let $x=\bigvee_{i=1}^{k}x_{i}$. Then $\text{max}\,B_{\leq x}=\{x_{1},\ldots,x_{k}\}$.
@@
De Concini and Procesi introduced three equivalent characterizations of a building set [14, Section 2.3]. Their third and most concise characterization has been featured prominently for the Boolean lattice in the work of Postnikov [32, Definition 7.1] and Feichtner-Sturmfels [21, Lemma 3.9] [^6] We note that building sets for the Boolean lattice have also appeared in the work of Schmitt under the name Whitney systems [33, Definition 1]. [^7] Building sets for the Boolean lattice correspond to certain smooth polytopes now known as nestohedra., but their characterization has been largely overlooked by researchers in matroid theory. Recently, Bibby-Denham-Feicthner proved one half of this characterization holds for locally geometric semilattices [7, Propositionn 2.5.3 (b)], a result which was utilized by Pagaria-Pezzoli [31]. Here we extend de Concini-Procesi’s characterization to general meet-semilattices and note that this appears to be the first time this characterization has been described in full for geometric lattices. For recovering the Boolean case from Proposition 2.11, note that here $I(\mathcal{L})=\{\{i\}:i\in[n]\}$, meets and joins are intersections and unions, respectively, and joins always exist.
@@
Let $\mathcal{L}$ be a meet-semilattice and $B\subseteq\mathcal{L}^{+}$. Then $B$ is a building set if and only if
@@
()($\Leftarrow$)( ⇐ ) Let $B\subseteq\mathcal{L}^{+}$ which satisfies conditions (1) and (2). Let $x\in\mathcal{L}^{+}$ and let $\text{max}\,B_{\leq x}=\{x_{1},...,x_{k}\}$. We wish to show $\prod_{j=1}^{k}[\hat{0},x_{j}]\cong[\hat{0},x]$. We begin with the observation that for each $1\leq i,j\leq k$ with $i\neq j$, $[\hat{0},x_{i}]\cap[\hat{0},x_{j}]=\hat{0}$. Suppose this is not the case and let $z\in[\hat{0},x_{i}]\cap[\hat{0},x_{j}]$ with $\hat{0}\neq z$. Then $x_{i}\wedge x_{j}\neq\hat{0}$ and $x_{i},x_{j}<x$, so $x_{i}\vee x_{j}\in\mathcal{L}$, thus $x_{i}\vee x_{j}\in B$. But $x_{i},x_{j}<x_{i}\vee x_{j}$ and $x_{i}\vee x_{j}\leq x$ which contradicts the maximality of $x_{i}$ and $x_{j}$.
Let $\text{max}\,I(\mathcal{L})_{\leq x}=\{y_{1},\ldots,y_{l}\}$. For each $1\leq t\leq k$ let $\pi_{t}=\{y_{i}\in\text{max}\,I(\mathcal{L})_{\leq x}:y_{i}\leq x_{t}\}$. We claim that the set of $\pi_{t}$ forms a partition of $\text{max}\,I(\mathcal{L})_{\leq x}$. By assumption, $I(\mathcal{L})$ is a subset of $B$, so for each $y_{i}$ there exists some $x_{t}$ such that $y_{i}\leq x_{t}$, and we know by disjointness of the $[\hat{0},x_{i}]$, that $x_{t}$ is unique in this regard – this establishes the partition.
@@
[…]

### v4

@@ Abstract.
Building sets were introduced in the study of wonderful compactifications of hyperplane arrangement complements and were later generalized to finite meet-semilattices. Convex geometries, the duals of antimatroids, offer a robust combinatorial abstraction of convexity. Supersolvable convex geometries and antimatroids appear in the study of poset closure operators, Coxeter groups, and matroid activities. We prove that the building sets on a finite meet-semilattice form a supersolvable convex geometry. As an application, we demonstrate that building sets and nested set complexes respect certain restrictions of finite meet-semilattices, unifying and extending results of several authors.
@@ 1. Introduction
De Concini and Procesi’s celebrated wonderful compactifications of a hyperplane arrangement complement proceed by blow ups of certain flats of the arrangement [de1995wonderful]. There are choices for which flats are blown up, and each set of choices is encoded by a building set. Feichtner and Kozlov [feichtner2004incidence] extended the notion of a building set from the lattice of flats of a hyperplane arrangement to arbitrary finite meet-semilattices [^1] Henceforth, all posets in this paper will be assumed to be finite., and observed that this generalization admits connections to other topics such as stellar subdivisions of fans.
@@
The Bergman fan of a matroid was introduced by Ardila-Klivans [ardila2006bergman] as a matroidal generalization of the tropicalization of a linear space. Feichtner-Sturmfels [feichtner2005matroid] showed that the nested set complex associated to a building set for the lattice of flats gives a unimodular triangulation of the Bergman fan. In seminal work of Adiprasito-Huh-Katz [adiprasito2018hodge], the Chow ring of a matroid, introduced earlier by Feichtner-Yuzvinsky [feichtner2004chow], was utilized for settling the Heron-Rota-Welsh conjecture on the log-concavity of the coefficients of the characteristic polynomial of a matroid. Although Feichtner-Yuzvinsky’s construction is defined for arbitrary building sets, Adiprasito-Huh-Katz found that the maximum building set was sufficient for their goals. [^2] While the Bergman fans of [adiprasito2018hodge] are not Bergman fans induced by building sets on the lattice of flats, we note a connection in footnote 3.18. Much of the subsequent work by researchers studying Chow rings of matroids has been focused on the maximum building set with a few important notable exceptions: the conormal fan of a matroid [^3] The biflats, viewed as a subset of $\mathcal{L}(M)^{op}\times\mathcal{L}(M^{\perp})$, are closed under taking joins and thus satisfy condition (2) of Proposition 2.11. While the biflats do not satisfy condition (1) of Proposition 2.11, the image of the biflats in $\Sigma_{M}\times\Sigma_{M^{\perp}}$ contains the image of the irreducibles. As the biflats are closed under taking joins, the would-be nested set complex associated to the biflats is the order complex, of which the conormal fan is a distinguished codimension-2 subcomplex. This observation will be elaborated upon in forthcoming work of the first author and Castillo. [ardila2023lagrangianC, ardila2023lagrangian, nathanson2023topology], the augmented Bergman fan of a matroid [braden2020singular, braden2022semi, eur2023stellahedral, liao2022stembridge, mastroeni2023chow], the Bergman fan of a polymatroid [crowley2022bergman, eur2023intersection, pagaria2023hodge] [^4] It follows from each of the works [amini2020hodge, ardila2023lagrangianC, pagaria2023hodge] that the Kähler package holds for the Chow ring of a matroid with respect to an arbitrary building set., and a few additional works [coron2022matroids, coron2025supersolvability, eur2025building].
@@
Perhaps surprisingly, there has been no work exploring the systematic construction of building sets for the lattice of flats of a matroid – this apparent gap in the literature was the genesis of this paper. Although our primary motivation is the study of matroids, we have found that our results hold for arbitrary meet-semilattices. In their original article, de Concini-Procesi provided three different equivalent descriptions of a building set [de1995wonderful]. Their most elegant characterization was popularized by Postnikov [postnikov2009permutohedra] and Feichtner-Sturfmels [feichtner2005matroid] for the Boolean lattice, but has largely been overlooked by researchers studying matroids. We begin by extending this characterization to a general meet-semilattice $\mathcal{L}$, e.g. a geometric lattice. We apply this result to demonstrate that the building sets for $\mathcal{L}$ form an intersection lattice. It follows that given an arbitrary subset $X\subseteq\mathcal{L}$, there exists a smallest building set $B$ which contains $X$, thus the building sets for $\mathcal{L}$ determine a closure operator; this generalizes a result of Feichtner-Sturmfels [feichtner2005matroid] from the Boolean lattice to general meet-semilattices.
@@
Convex geometries are a special class of closure systems introduced independently by Edelman and Jamison [edelman1980meet, edelman1985theory, jamison1982perspective] providing a combinatorial abstraction of convex hulls in finite point configurations and upper ideals in posets. A fundamental result of Björner [bjorner1983matroids] states that convex geometries are dual to antimatroids, which are an important class of greedoids. Supersolvable antimatroids and convex geometries were introduced implicitly in the work of Stanley [stanley1972supersolvable], and were developed in earnest by Armstrong [armstrong2009sorting]. They arise naturally in the study of poset closure operators [hawrylycz1993lattice], Coxeter groups [armstrong2009sorting], and matroid activities [gillespie2020convexity]. We prove that the building sets on a meet-semilattice form a supersolvable convex geometry.
@@
We demonstrate the utility of our results by proving that the building sets and nested set complexes behave well with respect to restrictions of meet-semilattices. This result unifies the work of Mantovani-Padrol-Pilaud on oriented matroids [mantovani2025facial] with Bergman fan constructions arising in the study of Hodge theory for matroids due to Braden-Huh-Matherne-Proudfoot-Wang [braden2022semi], Crowley-Huh-Larson-Simpson-Wang [crowley2022bergman], and Eur-Larson[eur2023intersection].
@@ 2. Semilattices and Building Sets
In this section we review the theory of posets, semilattices, and building sets. We prove Proposition 2.11 which extends a characterization of building sets due to de Concini and Procesi from intersection lattices of subspace arrangements to general meet-semilattices. The section concludes with Proposition 2.12 which states that the collection of building sets on a meet-semilattice ordered by inclusion forms an intersection lattice, generalizing an observation of Feichtner-Sturmfels from the Boolean lattice to general meet-semilattices.
@@
We write $x<y$ if $x\leq y$ and $x\neq y$, and we define $x>y$ similarly. We say that $y$ covers $x$ if $x<y$ and there is no element $z\in\mathcal{P}$ such that $x<y<z$. We define the subset $\mathcal{P}_{\leq x}$ to be $\{z\in\mathcal{P}:z\leq x\}$. For a subset $B$ of $\mathcal{P}$ with the induced order, and $x\in\mathcal{P}$, we define $B_{\leq x}=\{z\in B:z\leq x\}$, and $\text{max}\,B_{\leq x}$ as the set of maximal elements in $B_{\leq x}$ with respect to $\leq$. An interval $x,y]\subseteq\mathcal{P}$ is the set of elements $z\in\mathcal{P}$ such that $x\leq z\leq y$. A linear extension $_{L}$ of a partially ordered set $(\mathcal{P},\leq)$ is a total order of $\mathcal{P}$ such that if $x,y\in\mathcal{P}$ with $x<y$, then $x<_{L}y$. Let $(\mathcal{P}_{1},\leq_{1})$ and $(P_{2},\leq_{2})$ be posets. A function $f:\mathcal{P}_{1}\rightarrow\mathcal{P}_{2}$ is an order embedding if $x\leq_{1}y$ precisely when $f(x)\leq_{2}f(y)$. Note that an order embedding is necessarily injective. If $f$ is an inclusion map, we may write $\mathcal{P}_{1}\hookrightarrow\mathcal{P}_{2}$. An order embedding is an isomorphism if it is bijective. Let $\{(\mathcal{P}_{k},\leq_{k}):1\leq k\leq n\}$ be a collection of posets.
[…]
@@
[feichtner2004incidence, Proposition 2.1] Let $(\mathcal{P},\leq)$ be a partially ordered set with a minimum element $\hat{0}$ and a maximum element $x$. Let $\text{max}\,I(\mathcal{P})=\{y_{1},...,y_{l}\}$, then
@@
[…]
@@
Let $\mathcal{P}$ be a poset and $x,y\in\mathcal{P}$. The greatest lower bound of elements $x$ and $y$, if it exists, is the meet of $x$ and $y$, denoted $x\wedge y$. The least upper bound of $x$ and $y$, if it exists, is the join of $x$ and $y$, denoted $x\vee y$. A poset $\mathcal{L}$ is called a meet-semilattice if the meet of any two of its elements exists in $\mathcal{L}$. We define a join-semilattice similarly. If $\mathcal{L}$ is a meet-semilattice and $x,y\in\mathcal{L}$, we write $x\vee y\in\mathcal{L}$ to mean the statement “The join of $x$ and $y$ exists in $\mathcal{L}$.” A poset $\mathcal{L}$ is called a lattice if it is a meet-semilattice and a join-semilattice. For avoiding confusion we may denote $\wedge$ as $\wedge^{\mathcal{L}}$, and $\vee$ as $\vee^{\mathcal{L}}$. If $\mathcal{L}$ is a meet-semilattice, we define $\{$$\mathcal{L}^{+}:=\mathcal{L}\backslash\{\hat{0}\}$. Let $\mathcal{L}$ and $\mathcal{K}$ be meet-semilattices. A meet-semilattice order embedding $f:\mathcal{L}\rightarrow\mathcal{K}$ is an order embedding of $\mathcal{L}$ and $\mathcal{K}$ as posets such that for every $x,y\in\mathcal{L}$, we have $x\wedge^{\mathcal{L}}y=f(x)\wedge^{\mathcal{K}}f(y)$.
@@
Thus if $\mathcal{L}$ is a meet-semilattice, or a join-semilattice, and $x,y\in\mathcal{L}$ with $x\leq y$, the interval $x,y]\subseteq\mathcal{L}$ is itself a lattice. Let $\mathcal{E}$ be a set and $\mathbb{S}$ a lattice of subsets of $\mathcal{E}$ with meet given by intersection, then we refer to $\mathbb{S}$ as an intersection lattice. [^5] We warn the reader that this terminology, although natural, could be a source of confusion as the lattice of flats of a hyperplane arrangement form an intersection lattice, but this is the reverse of the intersection order of flats in an arrangement. In de Concini-Procesi [de1995wonderful], this confusion is avoided by working in the dual vector space.
@@
Let $\mathcal{L},\mathcal{K}$ be meet-semilattices, $f:\mathcal{L}\rightarrow\mathcal{K}$ a meet-semilattice order embedding, and $x_{1},\ldots,x_{k}\in\mathcal{L}$.
${\rm If}\bigvee_{1\leq i\leq k}^{\mathcal{L}}x_{i}\in\mathcal{L}\,\,\,\,\,\text{then}\,\,\,\,\bigvee_{1\leq i\leq k}^{\mathcal{K}}x_{i}\leq\bigvee_{1\leq i\leq k}^{\mathcal{L}}x_{i}\,.$
@@
Because $f$ is an order embedding, we can identify $\wedge^{\mathcal{K}}=\wedge^{\mathcal{L}}$ and denote this meet by $\wedge$. Let $\mathcal{S}=\{y\in\mathcal{K}:\forall\,i\in[k],\,x_{i}\leq y\}$ and $\mathcal{T}=\{y\in\mathcal{L}:\forall\,i\in[k],\,x_{i}\leq y\}$, then $\mathcal{S}\supseteq\mathcal{T}$. Hence
@@
[de1995wonderful, Section 2.3][feichtner2004incidence, Definition 2.2] Let $\mathcal{L}$ be a meet-semilattice, and $B\subseteq\mathcal{L}^{+}$. The set $B$ is a building set of $\mathcal{L}$ if for any $x\in\mathcal{L}^{+}$ and $\text{max}\,B_{\leq x}=\{x_{1},...,x_{k}\}$ there is an isomorphism of posets
@@
with $\phi_{x}(y_{1},...,y_{k})=y_{1}\vee\cdots\vee y_{k}$ for $y_{j}\in[\hat{0},x_{j}]$ with $1\leq j\leq k.$ We call $\text{max}\,B_{\leq x}$ the set of factors of $x$ in $B$. We denote $\mathbb{B}(\mathcal{L})=\{B\subseteq\mathcal{L}^{+}:B\penalty 10000\ \text{is a building set of}\penalty 10000\ \mathcal{L}\}.$
@@
Let $\mathcal{L}$ be a meet-semilattice and $B\in\mathbb{B}(\mathcal{L})$. A subset $N\subseteq B$ is a nested set if for every $x_{1},...,x_{t}\in N$ with $t\geq 2$, which are incomparable elements, $\B\bigvee_{i=1}^{t}x_{i}\in\mathcal{L}\backslash$ B. The nested set complex $\mathcal{N}(B)$ is the abstract simplicial complex consisting of all nested sets of $B$. For avoiding confusion, we may denote $\mathcal{N}(B)$ as $\mathcal{N}_{\mathcal{L}}(B)$.
@@
[de1995wonderful, Section 2.4][feichtner2004incidence, Proposition 2.8] Let $\mathcal{L}$ be a meet-semilattice, $B\in\mathbb{B}(\mathcal{L})$, and $\mathcal{N}(B)$ the associated nested set complex. Let $N\in\mathcal{N}(B)$ with $x_{1}\ldots,x_{k}\in N$ incomparable and let $x=\bigvee_{i=1}^{k}x_{i}$. Then $\text{max}\,B_{\leq x}=\{x_{1},\ldots,x_{k}\}$.
@@
De Concini and Procesi introduced three equivalent characterizations of a building set [de1995wonderful, Section 2.3]. Their third and most concise characterization has been featured prominently for the Boolean lattice in the work of Postnikov [postnikov2009permutohedra, Definition 7.1] and Feichtner-Sturmfels [feichtner2005matroid, Lemma 3.9] [^6] We note that building sets for the Boolean lattice have also appeared in the work of Schmitt under the name Whitney systems [schmitt1995hopf, Definition 1]. [^7] Building sets for the Boolean lattice correspond to certain smooth polytopes now known as nestohedra., but this characterization has been largely overlooked by researchers in matroid theory. Recently, Bibby-Denham-Feichtner proved that one half of this characterization holds for locally geometric semilattices [bibby2022leray, Propositionn 2.5.3 (b)], a result which was utilized by Pagaria-Pezzoli [pagaria2023hodge]. Here we extend de Concini-Procesi’s characterization to general meet-semilattices and note that this appears to be the first time this characterization has been described in full for geometric lattices. To recover the Boolean case from Proposition 2.11, note that here $I(\mathcal{L})=\{\{i\}:i\in[n]\}$, meets and joins are intersections and unions, respectively, and joins always exist.
@@
Let $\mathcal{L}$ be a meet-semilattice and $B$ a subset of $\mathcal{L}^{+}$. Then $B$ is a building set if and only if
@@
$(\Leftarrow)$ Let $B\subseteq\mathcal{L}^{+}$ which satisfies conditions (1) and (2). Let $x\in\mathcal{L}^{+}$ and let $\text{max}\,B_{\leq x}=\{x_{1},...,x_{k}\}$. We wish to show $\prod_{j=1}^{k}[\hat{0},x_{j}]\cong[\hat{0},x]$. We begin with the observation that for each $1\leq i,j\leq k$ with $i\neq j$, $\hat{0},x_{i}]\cap[\hat{0},x_{j}]=\hat{0}$. Suppose this is not the case and let $z\in[\hat{0},x_{i}]\cap[\hat{0},x_{j}]$ with $\hat{0}\neq z$. Then $x_{i}\wedge x_{j}\neq\hat{0}$ and $x_{i},x_{j}<x$, so $x_{i}\vee x_{j}\in\mathcal{L}$, thus $x_{i}\vee x_{j}\in B$. But $x_{i},x_{j}<x_{i}\vee x_{j}$ and $x_{i}\vee x_{j}\leq x$ which contradicts the maximality of $x_{i}$ and $x_{j}$.
Let $\text{max}\,I(\mathcal{L})_{\leq x}=\{y_{1},\ldots,y_{l}\}$. For each $1\leq t\leq k$ let $\pi_{t}=\{y_{i}:y_{i}\leq x_{t}\}$. We claim that the set of $\pi_{t}$ forms a partition of $\text{max}\,I(\mathcal{L})_{\leq x}$. By assumption, $I(\mathcal{L})$ is a subset of $B$, so for each $y_{i}$ there exists some $x_{t}$ such that $y_{i}\leq x_{t}$, and we know by disjointness of the $\hat{0},x_{i}]$, that $x_{t}$ is unique in this regard – this establishes the partition.
@@
$\phi_{x}:\prod_{j=1}^{l}[\hat{0},y_{j}]\xrightarrow{\cong}[\hat{0},x]\hskip 14.22636pt\text{and}\hskip 14.22636pt\phi_{x_{t}}:\prod_{y\in\pi_{t}}[\hat{0},y]\xrightarrow{\cong}[\hat{0},x_{t}].$